News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 13 October 1999 11.00am EST

During the last 24 hours Chandra completed the GTO observations of the Tapesium Cluster, A2256 and M31. The observatory also passed through the final (11th) eclipse of the season at 10:31AM 10/13.

Real-time commanding was completed to reset power system software counters that had incremented during the last eclipse and to perform a reset one of the ACIS front-end processor units. Commanding was also performed during today's eclipse to reset the power system software to set the proper values for the eclipse timers following their reset to default values during the counter reset performed 10/12/99. The ACIS FEP reset was conducted after an increased number of parity errors were seen in data from the unit over the last 2 days. Initial data following the FEP reset indicates nominal error rates and the further assessment will consider the possibility of an SEU as the cause.

Analysis of ACIS data following the last radiation zone passage indicates that no further degradation has occurred to the energy resolution of the front-side illuminated chips.

The spacecraft systems continued to operate nominally.





Return to Status Reports